### v1

# ComparisonQA: Evaluating Factuality Robustness of LLMs Through Knowledge Frequency Control and Uncertainty

Abstract
The rapid development of LLMs has sparked extensive research into their factual knowledge. Current works claim that LLMs fall short on questions requiring less frequent knowledge. However, their proof is incomplete since they only study the influence of entity frequency, which can not fully represent knowledge frequency. So we introduce ComparisonQA benchmark, containing 283K abstract questions, each instantiated by a pair of high-frequency and low-frequency entities. It ensures a controllable comparison because the difference of knowledge frequency between such a pair is only related to entity frequency. In addition, to avoid possible semantic shortcuts, which is a severe problem of current LLMs study, we design a two-round method for knowledge robustness measurement utilizing both correctness and uncertainty. Experiments reveal that LLMs exhibit particularly low robustness regarding low-frequency knowledge, and GPT-4o is even the worst under this measurement. Besides, we introduce an automatic method to filter out questions with low-quality and shortcuts to form ComparisonQA-Hard. We find that uncertainty effectively identifies such questions while maintaining the data size.  [^1] https://github.com/HKUST-KnowComp/ComparisonQA
ComparisonQA: Evaluating Factuality Robustness of LLMs Through Knowledge Frequency Control and Uncertainty
Qing Zong, Zhaowei Wang, Tianshi Zheng, Xiyu Ren, Yangqiu Song Department of Computer Science and Engineering, HKUST {qzong, zwanggy, tzhengad, xrenaf, yqsong}@cse.ust.hk
1 Introduction
The rapid advancement of large language models (LLMs) has promoted a lot of study on their factual knowledge and reasoning ability, and many factuality QA benchmarks Wei et al. (2024a, b); Hendrycks et al. (2021); Kwiatkowski et al. (2019) have been introduced accordingly.
Figure 1: An example from ComparisonQA
Sun et al. (2024) and Mallen et al. (2023) find that LLM’s performance drops as the frequency of the knowledge required to solve the question decreases. They compare LLM’s performance on questions about entities with different frequencies. However, we argue that such comparisons are incomplete and unfair since entity frequency can not fully represent the knowledge frequency. For example, consider the two questions: When was [Einstein] born and What was [Einstein]’s favorite fruit. Although both questions involve well-known entities, they have quite different difficulties, since Einstein’s birthday can be found in many places while his favorite fruit is rarely mentioned.
To tackle the issue, we introduce our ComparisonQA benchmark. Pairs of high-frequency and low-frequency entities share the same question, as shown in Figure 1. The shared abstract question, with a hypernym to represent the two specific entities, guarantees that the difference of the required knowledge between such a pair is only related to the entities. This allows for a controllable comparison between high-frequency and low-frequency entities. ComparisonQA is a large scale dataset containing 283K such question pairs. It is generated through an automatic pipeline from raw knowledge, ensuring both diversity and scalability. Through this, we completely prove that LLM performance is closely related to the frequency of the knowledge required.
In our benchmark, we use the multiple-choice format Hendrycks et al. (2021); Geva et al. (2021) for a more robust and accurate evaluation. But semantic shortcuts Geirhos et al. (2020) between questions and options may help LLMs to guess the answer, which is a common but severe problem recently. Thus, we design a two-round method to evaluate LLMs’ robust knowledge through correctness and uncertainty. During experiments, we found that LLMs have very low robustness, especially on questions requiring low-frequency knowledge, where the powerful GPT-4o even performed the worst among all LLMs we tested.
We also provide a subset called ComparisonQA-Hard for future study. It contains 81K difficult low-frequency questions with high quality and no semantic shortcuts. We propose an automatic method that also uses correctness and uncertainty to filter out questions with low-quality and shortcuts. Experiments show that uncertainty is more effective in filtering these questions while maintaining the size of the benchmark, compared to the recent method used in benchmarks like SimpleQA Wei et al. (2024a), which only collects questions adversarially against LLMs’ responses.
In summary, we have three main contributions: (1) [Resource] We introduce ComparisonQA benchmark, where a pair of entities share the same abstract question. It enables a more reasonable proof that LLMs perform worse when the required knowledge is less frequent. (§3)
(2) [Method] We propose a two-round method using correctness and uncertainty for LLMs’ robust knowledge measurement. [Finding] LLMs can not stand such a test, especially on the low-frequency knowledge, where GPT-4o even performs the worst among all tested LLMs. (§4)
(3) [Resource] We introduce ComparisonQA-Hard benchmark containing only hard and low-frequency questions of high quality and no shortcuts. [Finding] When doing the automatic filtering, uncertainty is more helpful to find questions with low quality and shortcuts. (§5)
2 Related Works
2.1 Benchmarking LLMs’ Factuality
The factuality evaluation of LLMs has recently attracted significant attention Wang et al. (2024c). Some factuality benchmarks require open-ended generation by LLMs, such as SimpleQA Wei et al. (2024a), FreshQA Vu et al. (2024), SelfAware Yin et al. (2023), CLR-Fact Zheng et al. (2024a), and HaluEval Li et al. (2023). Such evaluations either rely heavily on expert annotation, or utilize automatic answer matching that sacrifices evaluation accuracy Min et al. (2023); Chern et al. (2023); Wang et al. (2024b). Other benchmarks adopt the format of Yes-or-No questions Geva et al. (2021); Zhang et al. (2024) or multiple-choice questions (MCQ) Hendrycks et al. (2021); Wang et al. (2024a). These formats allow model responses to be easily parsed and compared with gold labels, enabling solid yet efficient evaluations.
Figure 2: An overview of our benchmark curation pipeline. It contains three parts. Through the first two parts, (1) Entity Pairs Extraction from DBpedia and (2) Abstract Question Generation, we can get the whole ComparisonQA. And through the third part, (3) Simple Low-Quality Question Filtering, we can get a harder subset, containing only difficult low-frequency questions with high quality and no semantic shortcut.
2.2 Long-Tail Knowledge
Long-tail knowledge Wei et al. (2024b); Chen et al. (2023) is an important aspect of factuality. Kumar et al. (2024) proposes an automatic approach to generate questions for tail entities. Kandpal et al. (2023) find LLMs struggle to learn long-tail knowledge. Other works study the influence of knowledge frequency: Mallen et al. (2023) introduced PopQA, a long-tail benchmark, and found that models’ performance will change with the frequency of entities in the questions. Sun et al. (2024) also proves this by constructing questions around head, torso, and tail entities. However, questions in these benchmarks are all in the form of open-ended generation, which can not be easily evaluated. They also depend on limited number of templates to produce QA questions from knowledge graphs, which will significantly harm the diversity of the benchmarks. More importantly, they only consider the frequency of the entities involved but ignore the entire required knowledge.
2.3 Abstraction Knowledge
Existing works have studied various aspects of abstraction, for example, entity abstraction Wu et al. (2012); Song et al. (2015), event abstraction Wang et al. (2024e, d), and conceptual abstraction Han et al. (2024). Abstraction has been proven beneficial to downstream tasks like commonsense reasoning, numerical reasoning, and logical reasoning Zhou et al. (2024); Hong et al. (2024). In this paper, we control question difficulty by sharing the same abstraction form between a pair of entities (§3.1).
3 ComparisonQA
To ensure more fair and controllable comparisons between LLMs’ performance on high-frequency and low-frequency factual knowledge, we propose a new benchmark called ComparisonQA.
3.1 Question Formulation
For accurate evaluation, our questions are in the form of multiple-choice. Although recently there are several generative QA benchmarks, like SimpleQA Wei et al. (2024a), and also some automatic methods Zheng et al. (2023) to evaluate generated answers. There are still significant limitations in such generative QAs since the answer should be single and indisputable. Questions like What is the primary focus or intention behind Civil Procedure Rules cannot be included since there are many ways to answer this question. However, multiple-choice questions do not have such limitations.
Each piece of data in ComparisonQA, shown in Figure 1, contains an abstract question shared by two entities having the same hypernym. One entity, having many relationships in DBpedia Auer et al. (2007), which will be introduced next, is the high-frequency entity, and the other, with only a few relationships, is the low-frequency one. The question will have different correct answers for the two entities, respectively. Such kind of data form can help us make more detailed and controllable comparisons, and that is why we call it ComparisonQA. The whole benchmark, containing 283K such question pairs, is constructed through a fully automated pipeline, which is cheap and scalable.
3.2 Curation Pipeline
In this section, we discuss the data curation pipeline for our dataset. As shown in Figure 2, the pipeline contains three parts: (1) Entity Pairs Extraction from DBpedia, (2) Abstract Question Generation, and (3) Simple Low-Quality Question Filtering. Here we will introduce the first two parts, used to build ComparisonQA, and leave the third part, used to build ComparisonQA-Hard, for §5.
3.2.1 Entity Pairs Extraction from DBpedia
Following Sun et al. (2024), an entity’s frequency is defined by its number of relationships in DBpedia. High-frequency and low-frequency entities, whose cumulative relationships account for the first and last $1/3$ of all sorted entities, are those with relationships above 185 and below 107 separately, with experiments shown in Appendix B.
In practice, we first get all the hypernyms in DBpedia and classify the entities belonging to them into high-frequent and low-frequent based on these two relationship counts. Then, we map the entities one-by-one to get entity pairs. The reason behind this is that not every pair of entities can have a shared abstract question easily. For example, it’s hard to write a shared question for Einstein and Apple even though they are all high frequency entities. But it’s easy for two specific universities to share a same abstract question. In order for our entity pairs to produce high quality abstract questions, we ensure that the two entities have the same hypernym. This allows them to have similar descriptions, making it easier to generate an abstract question.
3.2.2 Abstract Question Generation
After acquiring the entity pairs, we adopt a multi-step curation pipeline to generate high-quality abstract multiple-choice questions: (1) Generate an abstract question without options according to the DBpedia descriptions of paired entities. (2) Separately generate the corresponding answers based on the descriptions, with length control to alleviate bias among candidate answers. (3) Generate distractors for both entities with length control. Compared to randomly selected distractors, LLM-selected distractors generation have demonstrated its effectiveness for the high relevance between distractors and designated choices Zheng et al. (2024b). (4) Formulate the final multiple-choice question using the four answer candidates above. (5) Proofread the question according to the standards below.
The standards for questions are presented as follows: (1) Quality: The questions should have one and only one correct answer for both entities. (2) Semantic Shortcuts: The correct answer cannot be simply guessed by the names of the entities or the way the questions are asked. For example, the question: What was the primary operational location of Sydney O-Class Tram? A. Oslo, Norway B. Sydney, Australia C. Stockholm, Sweden D. Melbourne, Australia is not allowed since only the correct answer contains Sydney which is also in the entity’s name. (3) Length Bias: The four options in one question should have roughly the same length to avoid length bias. In summary, the generated questions should have high-quality and no shortcuts, in semantics or length.
During the curation, we utilize GPT-4o-mini OpenAI (2024a) with few-shot expert-written Chain-of-thought (CoT) demonstrations Wei et al. (2022), with details provided in Appendix C.
3.3 Expert Verification
We enlist the help of three postgraduate students, each with extensive experience in NLP research, to validate the quality of these generated questions through a sample of 200 question pairs. The quality of each pair is decided by majority voting, with standard shown above. Results show that their total agreement (all 3 experts agree with the quality) is 88.0%. And 95.5% of the abstract questions are considered correct and of high quality both for the high frequency entity and the low frequency entity, demonstrating the reliability of our benchmark.
3.4 Main Evaluations
ComparisonQA is a large-scale benchmark comprising a total of 283,455 abstract questions, each paired with a high-frequency and a low-frequency instantiation. Detailed statistics are in Appendix A.
We experiment with a selection of LLMs on our curated ComparisonQA benchmark to investigate their performance on high-frequency questions and low-frequency questions and also the difference between them.
Models High Freq Question Low Freq Question Average Difference (H –> T)
Unc. Acc Ma-F1 Unc. Acc Ma-F1 Unc. Acc Ma-F1 Unc. Acc Ma-F1
Random - 25.29 25.29 - 25.22 25.22 - 25.26 25.26 - ↓ 0.07 ↓ 0.07
Majority - 25.70 10.22 - 25.14 10.04 - 25.42 10.13 - ↓ 0.56 ↓ 0.18
LLM (Open Source) + Zero-Shot
Llama-3 8B 54.33 65.90 63.60 81.54 53.83 51.29 67.94 59.87 57.44 06.11 ↓ 12.07 ↓ 12.30
Llama-3-Instruct 8B 77.55 80.72 80.71 117.41 69.03 68.95 97.48 74.88 74.83 ↑ 39.86 ↓ 11.69 ↓ 11.76
Llama-3.1 8B 55.91 65.29 63.31 83.35 52.66 50.52 69.63 58.98 56.92 ↑ 27.44 ↓ 12.63 ↓ 12.78
Llama-3.1-Instruct 8B 58.97 80.06 80.08 87.05 69.99 69.94 73.01 75.03 75.01 ↑ 28.08 ↓ 10.07 ↓ 10.14
Gemma-2 9B 124.97 64.50 64.80 211.34 52.24 51.23 168.16 58.37 58.01 86.37 ↓ 12.26 13.57
Phi-3.5-mini-Instruct 4B 27.81 72.81 72.78 39.80 65.26 65.00 33.81 69.04 68.89 ↑ 11.99 07.55 07.78
Falcon2 11B 56.93 70.72 69.80 87.31 58.07 56.70 72.12 64.40 63.25 ↑ 30.38 12.65 ↓ 13.10
Mistral-v0.3 7B 39.83 65.55 63.11 56.99 53.36 50.26 48.41 59.46 56.69 ↑ 17.16 ↓ 12.19 ↓ 12.85
Mistral-v0.3-Instruct 7B 44.73 73.53 73.14 66.09 63.05 62.40 55.41 68.29 67.77 ↑ 21.36 ↓ 10.48 ↓ 10.74
LLM (Open Source) + Few-Shot
Llama-3 8B 21.89 75.57 75.55 23.75 61.00 61.01 22.82 68.29 68.28 ↑ 1.86 14.57 14.55
Llama-3-Instruct 8B 26.20 79.94 79.92 28.70 67.98 67.95 27.45 73.96 73.93 2.50 ↓ 11.96 ↓ 11.96
Llama-3.1 8B 20.82 74.91 74.89 22.62 62.00 62.00 21.72 68.46 68.45 ↑ 1.81 ↓ 12.91 ↓ 12.90
Llama-3.1-Instruct 8B 20.63 79.74 79.74 22.48 69.09 69.07 21.56 74.42 74.40 ↑ 1.85 ↓ 10.65 ↓ 10.66
Gemma-2 9B 20.26 80.10 80.08 22.36 68.36 68.31 21.31 74.23 74.20 ↑ 2.10 ↓ 11.74 ↓ 11.77
Phi-3.5-mini-Instruct 4B 11.02 73.68 73.67 11.85 67.46 67.33 11.44 70.57 70.50 ↑ 0.83 06.22 06.34
Falcon2 11B 16.42 77.11 77.01 17.77 65.92 65.75 17.10 71.52 71.38 ↑ 1.35 ↓ 11.19 ↓ 11.26
Mistral-v0.3 7B 13.89 75.55 75.53 14.99 62.88 62.85 14.44 69.22 69.19 ↑ 1.10 ↓ 12.67 ↓ 12.68
Mistral-v0.3-Instruct 7B 15.97 74.46 74.40 17.40 65.51 65.35 16.69 69.99 69.87 ↑ 1.43 08.95 09.05
LLM (Proprietary) API
GPT4o-mini (Zero-Shot) 13.52 85.61 85.58 18.34 73.85 73.73 15.93 79.73 79.66 04.82 ↓ 11.76 ↓ 11.85
GPT4o-mini (Few-Shot) 25.74 84.78 84.69 38.17 72.76 72.47 31.96 78.77 78.58 ↑ 12.43 12.02 12.22
GPT4o-mini (CoT) 10.53 86.25 86.25 12.27 74.39 74.40 11.40 80.32 80.32 01.74 ↓ 11.85 ↓ 11.85
GPT4o (Zero-Shot) 14.18 93.86 93.95 30.98 85.76 86.69 22.58 89.81 90.32 ↑ 16.80 08.10 07.26
GPT4o (Few-Shot) 28.41 93.94 93.95 45.81 86.54 86.75 37.11 90.24 90.35 17.40 07.40 07.20
GPT4o (CoT) 10.39 92.40 92.47 18.36 85.47 85.72 14.38 88.93 89.10 07.97 06.93 0 6.75
Table 1: Performance of various LLMs on the testing set of ComparisonQA. Unc., Acc, and Ma-F1 denote Uncertainty, Accuracy, and Macro F1-score. The Difference column shows how scores change from high-frequency questions to low-frequency questions. The best performances within each method are underlined, and the best among all methods are bold-faced. And for the Difference column, We underline the largest difference within each method and bold the one among all methods. More results can be seen in Table 7.
3.4.1 Experiment Setup
Metric: We calculate the Uncertainty, Accuracy, and Macro F1-score between predictions and ground truth labels to evaluate the models. The calculation of uncertainty is explained in Appendix D.
Models: We experiment with 16 different models, with a full list in Appendix E, and categorize the evaluation into three types: (1) Open Source LLM Zero-Shot Qin et al. (2023). (2) Open Source LLM Few-Shot Brown et al. (2020). (3) Proprietary LLM API.
3.4.2 Results and Analysis
Evaluation results are reported in Table 1. Our observations include: (1) Huge drop in performance from high-freq to low-freq: All models suffer a performance decrease from high-frequency questions to low-frequency questions in all three settings. For instance, the accuracy and Macro F1-score of Llama-3-8B drop up to about 14 points in the few-shot setting. Proprietary models like GPT-4o are no exception. These all prove that LLM’s performance is closely related to the frequency of the knowledge in the corpus. (2) Increased Uncertainty from high-freq to low-freq: Similarly, the uncertainty of LLMs all increase from high-frequency questions to low-frequency questions. For example, the uncertainty difference of Gemma-2-9B is up to about 86 points when using zero-shot. The uncertainty in the zero-shot setting is generally higher than in the few-shot setting, and the difference is also more pronounced. We think it may be because LLMs find some familiar examples in the few-shot setting, which decreases their uncertainty. In spite of this, the difference is still clear between high-freq and low-freq. These all prove that LLMs not only perform better but also are more confident about high-frequency knowledge. (3) Few-shot helps a lot only for LLMs without instruction-tuning: Most non-instruction-tuned LLMs show a huge improvement in performance from zero-shot to few-shot, but performance is similar for those instruction-tuned LLMs. This could be because the few-shot examples only teach LLMs how to do multiple-choice questions, while those instruction-tuned ones have already learned. (4) CoT lowers uncertainty but does not aid performance: It’s obvious that after the CoT inference, LLMs are more sure about their answers. However, results show that their performance does not have a big difference, which means CoT can not help with such factuality questions.
Models First Round Second Round
High Low Avg. Diff High Low Avg. Diff
Open Source LLM
Llama-3 8B 75.57 61.00 68.29 14.57 66.09 (-9.48)0 39.15 (-21.86) 52.62 (-15.67) ↓ 26.95 (+12.38)
Llama-3-Instruct 8B 79.94 67.98 73.96 ↓ 11.96 69.04 (-10.91) 43.50 (-24.49) 56.27 (-17.69) ↓ 25.54 (+13.58)
Llama-3.1 8B 74.91 62.00 68.46 ↓ 12.91 66.27 (-8.64)0 40.35 (-21.65) 53.31 (-15.15) ↓ 25.92 (+13.01)
Llama-3.1-Instruct 8B 79.74 69.09 74.42 ↓ 10.65 72.89 (-6.85)0 43.91 (-25.19) 58.40 (-16.02) 28.98 (+18.34)
Llama-3.2 3B 68.89 57.26 63.08 ↓ 11.63 59.15 (-9.73)0 38.19 (-19.07) 48.67 (-14.41) ↓ 20.96 (+9.33)0
Llama-3.2-Instruct 3B 71.43 62.12 66.78 09.32 64.14 (-7.30)0 40.19 (-21.93) 52.17 (-14.62) ↓ 22.68 (+14.63)
Gemma-2 2B 62.99 50.93 56.96 ↓ 12.06 52.54 (-10.44) 35.40 (-15.53) 43.97 (-12.99) ↓ 17.14 (+5.09)0
Gemma-2 9B 80.10 68.36 74.23 ↓ 11.74 71.26 (-8.84)0 46.98 (-21.38) 59.12 (-15.11) ↓ 24.28 (+12.54)
Phi-3.5-mini 4B 73.68 67.46 70.57 06.22 67.08 (-6.60)0 44.39 (-23.06) 55.74 (-14.84) ↓ 22.69 (+16.47)
Falcon2 11B 77.11 65.92 71.52 ↓ 11.19 64.27 (-12.84) 43.37 (-22.55) 53.82 (-17.70) ↓ 20.90 (+9.71)0
Mistral-v0.3 7B 75.55 62.88 69.22 ↓ 12.67 68.75 (-6.80)0 39.81 (-23.07) 54.28 (-14.94) ↓ 28.94 (+16.27)
Mistral-v0.3-Instruct 7B 74.46 65.51 69.99 08.95 68.25 (-6.21)0 41.19 (-24.32) 54.72 (-15.27) ↓ 27.06 (+18.10)
Proprietary LLM
GPT4o-mini 84.78 72.76 78.77 12.02 71.00 (-13.78) 33.16 (-39.60) 52.08 (-26.69) ↓ 37.83 (+25.81)
GPT4o 93.94 86.54 90.24 07.40 79.85 (-14.09) 25.50 (-61.04) 52.68 (-37.57) 54.34 (+46.94)
Table 2: Accuracy scores in LLMs’ robust knowledge measurement. We also report the changes in the scores from the first round to the second round. The best performances within each method are underlined, and the best among all methods are bold-faced. And for the Difference column and values in parentheses, We underline the largest difference within each method and bold the one among all methods.
4 Robust Knowledge Measurement
With the help of ComparisonQA, we can conduct a more detailed and controllable study of the factual knowledge of LLMs.
When considering how humans tackle multiple-choice questions, it’s often the case that we do not really know the correct answers. Instead, we rely on semantic shortcuts within the questions to make educated guesses. Although we intentionally exclude these shortcuts when constructing the benchmark, they are difficult to eliminate entirely from multiple-choice questions. Such a situation often occurs even in human exam questions. Therefore, we need to devise an effective method to evaluate the robustness of factual knowledge of LLMs in the form of multiple-choice questions.
4.1 The Definition of Robust Knowledge
We categorize LLM’s results into four scenarios based on its uncertainty and the correctness of its answers. The uncertainty pertains to the statement based on pure Question+Answer (without options), which isolates the correct knowledge without the influence of the three distractors. Details are in Appendix D. Meanwhile, the correctness refers to the original multiple-choice question.
Low Uncertainty & Correct Answer: LLM shows confidence about the correct knowledge and also answers the question correctly. In this case, we consider the LLM to possess robust knowledge.
High Uncertainty & Incorrect Answer: LLM expresses uncertainty and answers incorrectly. So, we conclude that the LLM lacks the knowledge.
High Uncertainty & Correct Answer: LLM is unsure but answers correctly. This indicates that it may retain the correct knowledge, but the memory is vague, or that the semantic shortcuts in the question lead to the correct answer.
Low Uncertainty & Incorrect Answer: LLM is confident yet answers incorrectly. This could result from the LLM recalling incorrect knowledge or from misleading distractors in questions.
In the first two categories, we can determine whether the LLM truly possesses the knowledge. However, in the latter two cases, multiple factors influence the final results, and the LLM’s grasp of knowledge is not robust.
This classification method leverages the strengths of both multiple-choice and generative questions, since we collect the uncertainty score without the distractors. While multiple-choice questions are easy to evaluate, they may allow for shortcuts; generative questions, on the other hand, are the opposite. Our method capitalizes on the uncertainty inherent in generative questions, which do not have shortcuts, and the accuracy of easily parsed answers provided by multiple-choice questions. This approach ensures that evaluation remains straightforward while fully addressing the potential for shortcuts.
Figure 3: Heatmaps illustrating how subset quality changes with incorrect model number and high uncertainty remaining ratio. The former refers to the minimum number of times that each remaining question in the subset is answered incorrectly. The latter refers to the proportion of high-uncertainty questions that are retained in the subset.
4.2 The Two-Round Measurement
We introduce our two-round measurement, which can be applied to any multiple-choice benchmarks, based on the four categories. In the first round, we present multiple-choice questions to evaluate LLM’s performance. Then, results are classified into four categories. For questions falling into the latter two cases (high-uncertainty correct & low-uncertainty incorrect), we will conduct a second round questioning. Scores will be modified if the correctness of any questions changes in this round.
For questions requiring reassessment in the second round, we ask LLMs to judge whether the four statements, with details in Appendix D, are true or false. The LLM is considered to truly possess the knowledge only when all four statements are accurately judged. On one hand, the second round provides an opportunity for LLMs to correct the answer by breaking distractors into separate questions, and on the other hand, it identifies questions where the LLMs simply guess the correct answers.
Compared to breaking down the multiple-choice questions into four correctness judgments directly from the start, our two-round approach offers a comprehensive analysis of how LLMs’ performance changes from the first to the second stage. It leverages both uncertainty and correctness, providing deeper insights into LLMs’ confidence and robustness regarding the factual knowledge they retain. Additionally, this method enables us to determine whether a particular result is due to LLM’s lack of knowledge or the shortcuts and misleading distractors in the multiple-choice questions.
Our ComparisonQA benchmark ensures that the classification and comparison is fair between high-frequency and low-frequency questions. Because we have a shared abstract question for each pair of entities, thus controlling other factors which may affect LLM’s uncertainty and accuracy. Thus, we can rule out all the other factors to purely compare the difference between high-frequency and low-frequency knowledge.
4.3 Experiments and Analysis
Here, we conduct the experiments using this method to measure LLMs’ knowledge robustness in a few-shot manner. Results are shown in Table 2
There are very interesting observations from the results: (1) GPT-4o and GPT-4o-mini can not stand the test of robust knowledge measurement especially on low-frequency knowledge: Their performance is still very good on high-frequency questions after the second round but drops a lot on low-frequency ones. The accuracy of GPT-4o on low-frequency questions after the second round is only 25.50, about 61 points lower than the first round. Its performance is even the worst of all the evaluated models, with GPT-4o-mini being the second-to-last. (2) The LLMs’ grasp of low-frequency knowledge is less robust than that of high-frequency knowledge: Performance of LLMs in the second round all drop much more on low-frequency questions than on high-frequency questions. For example, in all of the open source LLMs, the accuracy of Llama-3.1-8B-Instruct in the second round, whose average performance is the best in the first round, dropped up to about 25 points on low-frequency questions while only about 7 points on high-frequency questions. (3) LLMs which can stand the test of robust knowledge measurement must have a very accurate grasp of low-frequency knowledge: LLMs whose average accuracy in the second round is higher than 55 are all those who performed relatively better on the low-frequency questions. This illustrates the importance of the LLMs mastering low-frequency knowledge if they want to be reliable models in the aspect of factual knowledge.
5 ComparisonQA-Hard
Directly collected questions, especially the high-frequency part, are a bit simple for today’s LLMs. More importantly, they may have low quality and semantic shortcuts. So we introduce a subset called ComparisonQA-Hard for future study, containing only difficult, low-frequency questions that are filtered from ComparisonQA to ensure high quality and avoid semantic shortcuts.
5.1 Simple Low-Quality Question Filtering
The filtering process utilizes both correctness and uncertainty, similar to the last section. Previous benchmarks, like SimpleQA Wei et al. (2024a) and MMLU-Pro Wang et al. (2024a), were collected adversarially based on LLMs’ responses to ensure question difficulty. We enhance this method by also considering LLMs’ uncertainty to achieve the selection of high-quality, shortcut-free questions.
As illustrated in the third part of Figure 2, we collect the correctness and uncertainty from six open-source LLMs, listed in Appendix E, for questions with low-frequency entities. Different from above, both metrics here are about the entire multiple-choice questions, considering all the four options’ quality. We choose the questions that many models answer incorrectly and exhibit high uncertainty for our hard subset. We define two hyperparameters: incorrect model number and high uncertainty remaining ratio. The former is the minimum number of times that each remaining question is answered incorrectly. The latter is the proportion of high-uncertainty questions (sort by the sum of the uncertainty of all models) retained in the subset.
The method is designed based on this assumption: Low-quality questions and those with semantic shortcuts are often associated with correct answers and low uncertainty across different models, as our benchmark is constructed by LLMs-generated questions. The rationale behind this is as follows: (1) For low-quality questions, if the question or answer is incorrect, it is likely that the LLM did not generate it from the descriptions but rather from its internal knowledge. Consequently, models may display high confidence and yield correct answers. (2) For questions containing shortcuts, models may cheat through shortcuts, resulting in lower uncertainty and higher accuracy. These will all be proved in the following experiments.
Method Backbone Unc. Acc Ma-F1
Open Source LLM Llama-3 8B 25.04 27.95 27.74
Llama-3-Instruct 8B 30.57 22.38 22.32
Llama-3.1 8B 23.87 28.76 28.40
Llama-3.1-Instruct 8B 23.77 27.06 26.59
Llama-3.2 3B 26.97 27.10 27.05
Llama-3.2-Instruct 3B 27.04 28.68 28.65
Gemma-2 2B 31.07 22.80 22.35
Gemma-2 9B 23.67 24.71 24.64
Phi-3.5-mini 4B 12.42 31.87 31.19
Falcon2 11B 18.65 23.70 23.64
Mistral-v0.3 7B 15.69 23.96 23.74
Mistral-v0.3-Instruct 7B 18.36 23.65 23.25
Proprietary LLM GPT4o-mini 46.49 38.13 37.82
GPT4o 54.66 69.98 70.02
Table 3: Performance of various LLMs on the testing set of ComparisonQA-Hard. Unc., Acc, and Ma-F1, denote Uncertainty, Accuracy, and Macro F1-score. The best performances within each method are underlined and the best among all methods are bold-faced.
5.2 Parameters Chosen by Expert Verification
To validate our filtering method, we invite experts to annotate the shortcuts in the 200 randomly sampled questions in the same setting mentioned in §3.3. Results show that 9.4% of the 95.5% correct and high-quality questions are identified as having shortcuts which may decrease difficulty.
Then, we examine how the following metrics change with different settings regarding correctness and uncertainty, which are illustrated in Figure 3: (1) Remaining Low-Quality Ratio (the proportion of remaining low-quality problems relative to the original number of low-quality problems), (2) Remaining Shortcut Ratio (the proportion of remaining problems with shortcuts relative to the original number of problems with shortcuts), (3) GPT-4o Accuracy (the accuracy of GPT-4o on the remaining questions), and (4) Remaining Question Count (the size of the remaining questions). The results suggest that both uncertainty and correctness contribute to the selection of high-quality, shortcut-free questions, thereby demonstrating the effectiveness of our method. While it is expected that different models’ correctness would aid in identifying more challenging questions, it is noteworthy that uncertainty proved to be more effective in selecting high-quality and shortcut-free questions while maintaining the dataset size, validating the inclusion of uncertainty in our filtering method.
Finally, we choose to set incorrect model number to 3 and high uncertainty remaining ratio to 0.8 according to Figure 3. It is a trade-off between quality, difficulty, and subset size. In this setting, only 1.5% of the total questions are of low quality, and 2.1% with shortcuts. Finally, the subset size is 81K, with a GPT-4o accuracy of 70%. Detailed statistics are shown in Appendix A.
5.3 Experiments and Analysis
Then we conduct experiments on ComparisonQA-Hard in a few-shot manner, with results shown in Table 3. In the multiple-choice question format, the open-source LLMs all significantly underperform, indicating the difficulty of our benchmark. For the proprietary LLMs, GPT-4o-mini also has a poor performance with an accuracy of about 38, even though we do not use it when constructing the subset. GPT-4o is better with an accuracy of about 70, but still has a huge room for future enhancement. And, predictably, its knowledge robustness will drop much more on this benchmark according to our experiments in §4.3.
6 Conclusions
In this paper, we first introduce ComparisonQA benchmark to evaluate LLMs’ factual knowledge, with a fully automatic pipeline. This benchmark allows for more controllable and detailed comparisons between high-frequency and low-frequency knowledge of LLMs. Then, we propose a two-round method utilizing correctness and uncertainty to measure LLMs’ knowledge robustness. And we are surprised to find that GPT-4o can not stand such a test, especially on the low-frequency knowledge, and even performs the worst among all LLMs we test. At last, we provide a subset called ComparisonQA-Hard, which contains only difficult low-frequency questions of high quality and no shortcuts for future study. During the question filtering, we also discovered that uncertainty is more helpful in filtering out questions with low quality and shortcuts compared with correctness, which is often used by recent works.
Limitations
While we contribute valuable resources, methods, and findings to advance the probing of LLMs’ factual knowledge, several limitations still exist that cannot be covered in this single work.
In this paper, we provide a shared abstract question with the entities being the only varying part. This approach ensures that the difference in knowledge frequency between a pair is only related to entity frequency. Future research could investigate methods for measuring the entire knowledge frequency in training data, rather than limiting it to entity frequency. We believe it is a challenging but valuable task. Additionally, our focus is on fixed factual knowledge. But those fast-changing knowledge also deserves attention. Specifically, exploring how knowledge frequency can assist LLMs in acquiring new information is a worthwhile area for further study.
Additionally, our focus is on fixed knowledge, but fast-changing factual knowledge Do et al. (2024b) and other knowledge Do et al. (2024a); Wang et al. (2022, 2023) also deserves attention. Specifically, exploring how knowledge frequency can assist LLMs in acquiring new information Choi et al. (2023); Zong et al. (2023) is a worthwhile area for further study.
Ethics Statement
Offensive Content Elimination. Our benchmark curation pipeline, which involves generating content using LLMs, requires stringent measures to ensure that generated responses are free from offensive material. We manually review a random sample of 200 data instances from ComparisonQA for any offensive content. Based on our annotations, we have not detected any offensive content. Therefore, we believe our dataset is safe and will not yield any negative societal impact.
Licenses. We will share our code under the MIT license, allowing other researchers free access to our resources for research purposes. Our dataset will be released under a CC license, also providing scholars with free access. We take full responsibility for any rights violations or issues related to the data license. The DBpedia dataset used in this paper is shared under the CC BY-SA license, permitting its use for research. As for language models, we access all open-source LMs via the Huggingface Hub Wolf et al. (2020). All associated licenses permit user access for research purposes, and we have agreed to adhere to all terms of use.
Annotations. For expert verifications, we have obtained IRB approval and support from our institution’s department, enabling us to invite expert graduate students to validate the quality of our data. They all agree to participate voluntarily without being compensated. We have made significant efforts to eliminate offensive content, thereby ensuring that no annotators are offended.
Appendices
Appendix A Benchmark Statistics
ComparisonQA is a large-scale benchmark comprising a total of 283,455 abstract question pairs, each paired with a high-frequency and a low-frequency entity. We guarantee that each entity corresponds to a single question sourced from 9166 hypernyms, to ensure no overlap. We partition our data into training, validation, and testing splits following an 8:1:1 ratio, ensuring that entities of different frequency intervals are evenly distributed in each split. Specific details are shown in Table 4.
And for ComparisonQA-Hard, there are 81,136 high-quality and shortcut-free questions with low-frequency entities from 4,876 hypernyms in total, with details shown in Table 5.
Split #Q. Pair #Entity #Hyper. #H. R. #L. R.
Train 226,762 453,524 8,430 628 76
Valid 28,345 56,690 3,316 627 76
Test 28,348 56,696 3,314 629 76
Total 283,455 566,910 9,166 628 76
Table 4: The statistics of ComparisonQA benchmark. #Q. Pair refers to the number of question pairs. Hyper. means hypernym. #H. R. and #L. R. refer to the average number of relationships of high frequency entities and low frequency entities respectively.
Split #Q. = #Entity #Hyper. #L. R.
Train 65,057 4,396 77.45
Valid 7,978 1,466 77.23
Test 8,101 1,484 77.65
Total 81,136 4,876 77.44
Table 5: The statistics of ComparisonQA-Hard benchmark. Here the number of questions is equal to the number of entities since it only has low frequency entities.
Method Prompt
Zero-Shot [Question]
A. [OptionA]. B. [OptionB].
C. [OptionC]. D. [OptionD].
The correct answer is:
Few-Shot [Examples]
Answer the multiple choice question. Select only one correct answer from the choices, following above examples.
[Question]
A. [OptionA]. B. [OptionB].
C. [OptionC]. D. [OptionD].
CoT Question: [question]
Rational: [rationale]
Answer: **[option].**
Answer the multiple choice question. Think step by step and generate a short rationale to support your reasoning. Choose one best answer based on the generated rational, following the above format. Keep your whole response in 50 tokens.
[Question]
A. [OptionA]. B. [OptionB].
C. [OptionC]. D. [OptionD].
Table 6: The prompt used when evaluating LLMs on our benchmark. Placeholders [Examples], [Question], [OptionA], [OptionB], [OptionC], [OptionD] will be replaced with the real examples, questions and their options accordingly.
Appendix B Definition of High and Low Frequency
We randomly sample 1K entities from DBpedia and compute their relationship count. Then we sort these entities in order of their relationship count from highest to lowest. To make contrast significant, entities whose cumulative relationships account for the first $1/3$ of the total number of relationships are defined as the high frequency entities, and those accounting for the last $1/3$ are low frequency entities. Repeating this process 3 times, we get the number of relationships to define high frequency and low frequency entities, which are higher than 185 and lower than 107 separately. Of the 1K randomly sampled entities, 119 entities are classified as high frequent and 621 entities as low frequent.
Appendix C Details in ComparisonQA Construction
In the process of entity pairs extraction, hypernyms that do not have enough entities and low-frequency entities that do not have enough high-frequency entities to pair with are all discarded. Finally, we get 293K entities pairs from 9,261 hypernyms in total. These pairs are then fed into the LLM to generate questions. After LLM’s proofread stage, there are 283K question pairs from 9,166 hypernyms left, which are used to build our ComparisonQA benchmark.
In questions generated by LLMs, the answer for the high frequency entity is $A$, low frequency entity is $B$, and their distractors are $C$ and $D$ respectively. And in the end, we randomly shuffle the 4 options in one question to guarantee the balance of the correct options.
Appendix D Calculation of Uncertainty
Following Liu et al. (2023), we first combine the questions with each of their options, and use GPT-4o-mini to transform them into four statements. The data can also be found in our benchmark.
Then we compute the uncertainty for each statement. For open-source LLMs, uncertainty refers to their perplexity for generating the correct statement. For proprietary LLMs, we allow them to generate their uncertainty scores, as their perplexity is not accessible. The threshold between high and low uncertainty for each model is determined by its own average uncertainty on the benchmark respectively.
Appendix E Experiment Details
Models High Freq Question Low Freq Question Average Difference (H –> T)
Unc. Acc Ma-F1 Unc. Acc Ma-F1 Unc. Acc Ma-F1 Unc. Acc Ma-F1
Random - 25.29 25.29 - 25.22 25.22 - 25.26 25.26 - ↓ 0.07 ↓ 0.07
Majority - 25.70 10.22 - 25.14 10.04 - 25.42 10.13 - ↓ 0.56 ↓ 0.18
LLM (Open Source) + Zero-shot
Llama-3 8B 54.33 65.90 63.60 81.54 53.83 51.29 67.94 59.87 57.44 ↑ 6.11 ↓ 12.07 ↓ 12.30
Llama-3-Instr 8B 77.55 80.72 80.71 117.41 69.03 68.95 97.48 74.88 74.83 ↑ 39.86 ↓ 11.69 ↓ 11.76
Llama-3.1 8B 55.91 65.29 63.31 83.35 52.66 50.52 69.63 58.98 56.92 ↑ 27.44 ↓ 12.63 ↓ 12.78
Llama-3.1-Instr 8B 58.97 80.06 80.08 87.05 69.99 69.94 73.01 75.03 75.01 ↑ 28.08 ↓ 10.07 ↓ 10.14
Llama-3.2 3B 67.51 57.93 53.65 99.55 48.10 44.66 83.53 53.02 49.16 ↑ 32.04 ↓ 9.83 ↓ 9.00
Llama-3.2-Instr 3B 74.27 71.47 71.51 108.24 62.17 62.18 91.26 66.82 66.85 ↑ 33.97 ↓ 9.30 ↓ 9.33
Gemma-2 2B 133.57 46.48 43.40 213.40 37.76 34.07 173.49 42.12 38.74 ↑ 79.83 ↓ 8.72 ↓ 9.33
Gemma-2 9B 124.97 64.50 64.80 211.34 52.24 51.23 168.16 58.37 58.01 86.37 ↓ 12.26 13.57
Phi-3.5-mini-Instr 4B 27.81 72.81 72.78 39.80 65.26 65.00 33.81 69.04 68.89 ↑ 11.99 ↓ 7.55 ↓ 7.78
Falcon 7B 62.33 18.19 18.55 92.92 17.93 18.19 77.63 18.06 18.37 ↑ 30.59 ↓ 0.26 ↓ 0.36
Falcon-Instr 7B 88.88 25.68 14.49 128.75 25.71 14.01 108.82 25.70 14.25 ↑ 39.87 ↑ 0.02 ↓ 0.48
Falcon2 11B 56.93 70.72 69.80 87.31 58.07 56.70 72.12 64.40 63.25 ↑ 30.38 12.65 ↓ 13.10
Mistral-v0.3 7B 39.83 65.55 63.11 56.99 53.36 50.26 48.41 59.46 56.69 ↑ 17.16 ↓ 12.19 ↓ 12.85
Mistral-v0.3-Instr 7B 44.73 73.53 73.14 66.09 63.05 62.40 55.41 68.29 67.77 ↑ 21.36 ↓ 10.48 ↓ 10.74
LLM (Open Source) + 4-shot
Llama-3 8B 21.89 75.57 75.55 23.75 61.00 61.01 22.82 68.29 68.28 ↑ 1.86 14.57 14.55
Llama-3-Instr 8B 26.20 79.94 79.92 28.70 67.98 67.95 27.45 73.96 73.93 2.50 ↓ 11.96 ↓ 11.96
Llama-3.1 8B 20.82 74.91 74.89 22.62 62.00 62.00 21.72 68.46 68.45 ↑ 1.81 ↓ 12.91 ↓ 12.90
Llama-3.1-Instr 8B 20.63 79.74 79.74 22.48 69.09 69.07 21.56 74.42 74.40 ↑ 1.85 ↓ 10.65 ↓ 10.66
Llama-3.2 3B 23.58 68.89 68.84 25.59 57.26 57.17 24.59 63.08 63.01 ↑ 2.01 ↓ 11.63 ↓ 11.67
Llama-3.2-Instr 3B 23.66 71.43 71.43 25.67 62.12 62.09 24.67 66.78 66.76 ↑ 2.01 ↓ 9.32 ↓ 9.34
Gemma-2 2B 26.96 62.99 62.93 29.43 50.93 50.91 28.20 56.96 56.92 ↑ 2.47 ↓ 12.06 ↓ 12.02
Gemma-2 9B 20.26 80.10 80.08 22.36 68.36 68.31 21.31 74.23 74.20 ↑ 2.10 ↓ 11.74 ↓ 11.77
Phi-3.5-mini-Instr 4B 11.02 73.68 73.67 11.85 67.46 67.33 11.44 70.57 70.50 ↑ 0.83 ↓ 6.22 ↓ 6.34
Falcon 7B 21.69 28.40 21.25 23.42 28.25 20.63 22.56 28.33 20.94 ↑ 1.73 ↓ 0.15 ↓ 0.62
Falcon-Instr 7B 21.69 26.27 18.32 23.42 25.67 18.15 22.56 25.97 18.23 ↑ 1.73 ↓ 0.59 ↓ 0.17
Falcon2 11B 16.42 77.11 77.01 17.77 65.92 65.75 17.10 71.52 71.38 ↑ 1.35 ↓ 11.19 ↓ 11.26
Mistral-v0.3 7B 13.89 75.55 75.53 14.99 62.88 62.85 14.44 69.22 69.19 ↑ 1.10 ↓ 12.67 ↓ 12.68
Mistral-v0.3-Instr 7B 15.97 74.46 74.40 17.40 65.51 65.35 16.69 69.99 69.87 ↑ 1.43 ↓ 8.95 ↓ 9.05
LLM (Proprietary) API
GPT4o-mini (Zero-Shot) 13.52 85.61 85.58 18.34 73.85 73.73 15.93 79.73 79.66 ↑ 4.82 ↓ 11.76 ↓ 11.85
GPT4o-mini (Few-Shot) 25.74 84.78 84.69 38.17 72.76 72.47 31.96 78.77 78.58 ↑ 12.43 12.02 12.22
GPT4o-mini (CoT) 10.53 86.25 86.25 12.27 74.39 74.40 11.40 80.32 80.32 ↑ 1.74 ↓ 11.85 ↓ 11.85
GPT4o (Zero-Shot) 14.18 93.86 93.95 30.98 85.76 86.69 22.58 89.81 90.32 ↑ 16.80 ↓ 8.10 ↓ 7.26
GPT4o (Few-Shot) 28.41 93.94 93.95 45.81 86.54 86.75 37.11 90.24 90.35 17.40 ↓ 7.40 ↓ 7.20
GPT4o (CoT) 10.39 92.40 92.47 18.36 85.47 85.72 14.38 88.93 89.10 ↑ 7.97 ↓ 6.93 ↓ 6.75
Table 7: Performance of various LLMs on the testing set of ComparisonQA. Unc., Acc, and Ma-F1, denote Uncertainty, Accuracy, and Macro F1-score. And the Difference column shows how scores change from high frequency questions to low frequency questions. The best performances within each method are underlined and the best among all methods are bold-faced. And for the Difference column, We underline the largest difference within each method and bold the one among all methods.
For the main evaluations on ComparisonQA, we categorize the evaluation of different models into three types: (1) Open Source LLM Zero-Shot: We first evaluate Llama3, Llama3.1, Llama3.2 Touvron et al. (2023); Dubey et al. (2024), Gemma2 Mesnard et al. (2024); Riviere et al. (2024), Phi3.5 Abdin et al. (2024), Falcon, Falcon2 Malartic et al. (2024), Mistral Jiang et al. (2023), and their instruction versions accordingly in a zero-shot manner Qin et al. (2023). (2) Open Source LLM Few-Shot: Then we evaluate the above models in a few-shot manner Brown et al. (2020). Since our benchmark is in the form of four options multiple-choice questions, the shot number is set to four. (3) Proprietary LLM API: Finally, we evaluate the performance of GPT-4o OpenAI (2023, 2024b) and GPT-4o-mini OpenAI (2024a), using zero-shot, few-shot, and Chain-of-Thought (CoT; Wei et al., 2022). All the evaluation results are reported in Table 7.
And for the question filtering of ComparisonQA-Hard, the six open-source LLMs we use are Llama-3-8B-Instruct, Llama-3.1-8B-Instruct, gemma-2-9b, Phi-3.5-mini-instruct, Falcon-11B, and Mistral-7B-Instruct-v0.3.
All the open-source models are run on 4 NVIDIA A6000 (40G) GPUs with BF32. And for proprietary LLM, we access them via OpenAI API  [^2] https://platform.openai.com/docs/api-reference. The different kinds of prompts we use are shown in Table 6.